# Vision-Language Alignment from Compressed Image Representations using 2D Gaussian Splatting

### v1

@@ 6 Conclusion
    Joint representation–alignment training with novel model architectures better suited for 2DGS inputs;
@@
    Video and streaming extensions that model and learn from splats across time; and
@@
    End-to-end edge–cloud evaluation, measuring not only accuracy but also energy and latency benefits of transmission-efficient representations.
We position this work as a first step toward representation-first multimodal learning, where compact, structured inputs such as 2DGS become native substrates for alignment, moving beyond post-hoc pruning toward fundamentally more efficient and adaptive vision–language systems.
Acknowledgments

### v2

@@ 11.5 Model Architecture and Training Recipe Studies
    Rendering Back Into Pixels. We render our GS-1600 dataset back into RGB pixel space and retrain a ViT-B/16 (Small) encoder on the reconstructed images (Fig. 14). As expected for a lossy compressor, accuracy declines relative to training on original RGB, but the relative drop closely mirrors the performance drop observed when training CLIP using GS representations directly. This alignment suggests that our Perceiver-based GS encoder is extracting the majority of the semantic content available in the rendered images, and that the overall performance ceiling is fundamentally constrained by the quality of the compressed RGB surrogate. Because our GS encoder is distilled from an RGB-pretrained teacher, its achievable zero-shot performance is similarly bounded by the representational limits of the underlying pixel-domain ViT.
@@
    Training GS Encoders Fully From Scratch. We train our GS-1600 (196-token) encoder entirely from scratch, without RGB initialization, pretraining, or distillation (Fig. 15). This model exhibits substantially lower zero-shot accuracy and slower convergence, highlighting the absence of strong native inductive biases for 2DGS representations when trained without external guidance. In contrast, distillation from RGB features provides a powerful initialization signal that lifts the GS encoder out of the suboptimal optimization basin associated with scratch training, emphasizing the importance of cross-domain supervision for early-stage 2DGS models.
@@
    2DGS as a Tokenizer for Vanilla Transformers. To test the representational power of 2DGS in extreme compression regimes without specialized architectures, we feed a low number of Gaussian splats directly into a vanilla Transformer after lightweight per-channel normalization, rescaling, and Fourier feature augmentation (Fig. 16). We pre-train the GS encoder from scratch in this case. We evaluate extreme compression settings (196 points) and low ones (400 points). Accuracy lags behind architectures that explicitly map from $N$ splats to $M$ latents (e.g., Perceiver stems). The results indicate that starting from a richer Gaussian budget yields better fidelity and that an architectural bottleneck is required to appropriately condense spatially distributed Gaussians into semantically meaningful tokens. This points toward the need for future architectures that can adaptively produce the number of output tokens based on the density and structure of the input splats.
Figure 20: Rendered ViT: Zero-shot accuracy (left) and train loss (right).
Figure 21: Training From Scratch: Zero-shot accuracy (left) and train loss (right).
Figure 22: Token Ablation Study: Zero-shot accuracy (left) and train loss (right).